Your brother has a two-year old baby girl. Thinking ahead, he is considering saving for his daughter’s college education, which will begin in 16 years.

If the bank promises to pay your brother an interest rate of 4% per year, how much money will your brother need to put into his saving account today to have $2,000,000 in 16 years?


Show all your calculations.


1
Expert's answer
2020-11-03T11:27:31-0500

2000000=P(1+"\frac{4}{100}" )"^{14}"

2000000=P(1.04)"^{14}"

2000000=1.731856P

P="\frac{2000000}{1.731856}"

money he should save (P)=$1,154,830.40

Reason for using 14. The child is 2years now n she will be in college in 16years . Therefore the father can only save for her for 14years . 14 years to come the will be 16years